The division – wide aptitude test in Mathematics was conducted to students. The mean of the test is 58 and the standard deviation is 12. The scores also approximate the normal distribution. What percent of the scores is between 55 and 65? *


1
Expert's answer
2022-04-12T12:38:28-0400

"P(55<X<65)=P(\frac{55-58}{12}<Z<\frac{65-58}{12})=P(-0.25<Z<0.58)=0.71904-0.40129=0.31775"